OpenStudy (anonymous):

simplify the rational expression x2-3x+2 over x2+5x-6

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[\frac{x^2-3x+2}{x^2+5x-6}=\frac{(x-1)(x-2)}{(x-1)(x+6)}=\frac{x-2}{x+6}\]